Lillian Fok is president of Lakefront​ Manufacturing, a producer of bicycle tires. Fok makes 1000 tires per day with the following​ resources:                                                                                                     ​

Labor: 425 hours per day ​@​ $12.00 per hour

Raw​ Material: 21,000 pounds per day​ @ ​$1.00 per pound

​Energy: ​$5250 per day

​Capital: ​$10000 per day

c) The percent change in multifactor productivity if Fok can reduce the energy bill by ​$1000 per day without cutting production or changing any other inputs​ = ___ ​% ​(enter your response as a percentage rounded to two decimal​ places).
Answer #1

OLD PRODUCTIVITY

VALUES:
LABOR = 5100
ENERGY = 5250
MATERIALS = 21000
CAPITAL = 10000

TOTAL OUTPUT = 1000

PRODUCTIVITY = TOTAL OUTPUT / TOTAL INPUT

PRODUCTIVITY = 1000 / (5100 + 5250 + 21000 + 10000) = 0.0242


2. NEW PRODUCTIVITY

VALUES:
LABOR = 5100
ENERGY = 4250
MATERIALS = 21000
CAPITAL = 10000

TOTAL OUTPUT = 1000

PRODUCTIVITY = TOTAL OUTPUT / TOTAL INPUT

PRODUCTIVITY = 1000 / (5100 + 4250 + 21000 + 10000) = 0.0248


3. CHANGE IN PRODUCTIVITY
CHANGE IN PRODUCTIVITY = NEW PRODUCTIVITY - OLD PRODUCTIVITY / OLD PRODUCTIVITY * 100


((0.0248 - 0.0242) / 0.0242) * 100 = 2.48%


**I have rounded the intermediate calculations to 4 decimal places, if the answer is a little different, it might be due to the rounding off factor, please let me know if you have a specific rounding factor for the intermediate calculations and i will edit the answer.

**PLEASE LEAVE A POSITIVE RATING